40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 33     Entry time: Fri Jul 8 15:26:37 2016
Author: Alena 
Type: Facility 
Category: Daily Progress 
Subject: Vacuum chamber tests 

Fixed the 307 gauge controller (a missing contact on the rair panel). The low pressure gauge was connected to 1G port and has measured 1.7E-6 torr. We are not sure since how long the turbo was operating (no vacuum logger yet).

 

 

Attachment 1: WP_20160708_12_25_35_Pro.jpg  381 kB  | Hide | Hide all
WP_20160708_12_25_35_Pro.jpg
Attachment 2: WP_20160708_12_25_41_Pro.jpg  287 kB  | Hide | Hide all
WP_20160708_12_25_41_Pro.jpg
ELOG V3.1.3-